OpenStudy (anonymous):

Write an equation in point-slope form for this line through point J(4,1) with slope -4. Can you show me step by step so i can do other problems just like this.... thanks

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

use this \[\large y - y_1 = m(x-x_1)\] plug in m = -4, and the point (4, 1) is your (x1, y1)

OpenStudy (agent0smith):

Once you do that, you're done.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Which one?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

@agent0smith ^

OpenStudy (agent0smith):

Not really sure about 22, that's pretty hard to explain. You could draw two parallel lines of length AB, then make two perpendicular lines of length AB, and draw all four into a square. For the next, find the slope through each pair of points, using \[\large m = \frac{ y_2 - y_1}{ x_2 -x_1 }\] Once you get the two slopes, multiply them together. If it equals -1, they're perpendicular.

OpenStudy (anonymous):

can you plug them in for me

OpenStudy (anonymous):

for #23 mr @agent0smith

OpenStudy (agent0smith):

\[\large m = \frac{ -12 - (-10)}{ -5 -(-8) }\]and \[\large m = \frac{ -5 - (-6)}{ 17 -9 }\]